# Fourier Circuits in Neural Networks and Transformers: A Case Study of Modular Arithmetic with Multiple Inputs

### v3

@@ 5 EXPERIMENTS
Grokking under Different $k$. To support our data setting importance, we study grokking phenomenon on our data distribution. Following the experiments’ protocol in [69], we show there is the grokking phenomenon under different $k$. We train two-layer transformers with $m=160$ attention heads and hidden dimension as 128 on $k=2,3,4,5$-sum mod-$p=97,31,11,5$ addition dataset with 50% of the data in training. We use different $p$ to guarantee the dataset sizes are roughly equal to each other. Figure 4 shows that the grokking weakens as the number of $k$ increases, which is consistent with our analysis. When $k$ increases, the function class will become more complicated, as we may need more neurons to achieve the max-margin solution. Thus, we use our Theorem 4.1 as a metric to measure the data complexity. It implies that when the ground-truth function class becomes “complicated”, the transformers need to train more steps to fit the training datasets, and the generalization tends to be better. Brilliant recent works by [49, 45] argue that, during learning, the network will be first in the lazy training/NTK regime and then transfer to the rich/feature learning regime sharply, leading to a grokking phenomenon. We use learning steps required for regime switch as a metric of grokking strength.
@@
NTK is a notorious overparameterized regime, which probably needs a much larger number of neurons than our max-margin convergence case, i.e., much larger than $\Omega(2^{2k})$ in Theorem 4.1. Thus, under the fixed $m$ and increasing $k$, the model may easily escape the NTK regime, or there is no longer an NTK regime. Thus, we will see a weaker grokking phenomenon as the learning steps needed to transfer from the NTK regime to the feature learning regime become fewer. With increasing $k$, the model will have an “underfitting” issue in the NTK regime, meaning the model must need feature learning to fit the task but cannot only fit the task by NTK. However, the model still has an “overfitting” in the feature learning regime.
@@ 6 DISCUSSION
Grokking in Transformers. The interpretability of grokking in Transformers is explored in [63]. By examining various intermediate states within the residual stream of the Transformer model, it is validated that the model employs Fourier features to tackle the modular addition task. However, fully comprehending how the Transformer model and LLMs perform modular addition remains challenging based on the current work, particularly from a theoretical standpoint. We contend that beginning with a simplistic model setup and achieving a thorough and theoretical understanding of how the network utilizes Fourier features to address the problem serves as a valuable starting point and it provides a theoretical understanding of the grokking phenomenon. We believe that further study on Transformers will be an interesting and important future direction.
@@
Grokking, Benign Overfitting, and Implicit Bias. Recently, [103] connects the grokking phenomenon to benign overfitting [9, 13, 90, 25, 27]. It shows how the network undergoes a grokking period from catastrophic to benign overfitting. [49, 45] uses implicit bias [75, 29, 41, 83, 62, 12, 53, 38, 101, 102] to explain grokking, where grokking happens if the early phase bias implies an overfitting solution while late phase bias implies a generalizable solution. The intuition from the benign overfitting and the implicit bias well align with our observation in Section 5. It is interesting and valuable to rigorously analyze the grokking or emergent ability under different function class complexities, e.g., Eq (1). We leave this challenge problem as a future work.
Connection to Parity and SQ Hardness. If we let $p=2$, then $(a_{1}+\dots+a_{k})\bmod p$ will degenerate to parity function, i.e., $b_{1},\dots,b_{k}\in\{\pm 1\}$ and determining $\prod_{i=1}^{k}b_{i}$. Parity functions serve as a fundamental set of learning challenges in computational learning theory, often used to demonstrate computational obstacles [80]. In particular, $(n,k)$-sparse parity problem is notorious hard to learn, i.e., Statistical Query (SQ) hardness [8]. [22] showed that one-hidden layer networks need an $\Omega(\exp(k))$ number of neurons or an $\Omega(\exp(k))$ number of training steps to successfully learn it by SGD. In our work, we are studying Eq. (2), which is a more general function than parity and indeed is a learning hardness. Our Theorem 4.1 states that we need $\Omega(\exp(k))$ number of neurons to represent the maximum-margin solution, which well aligns with existing works. Our experiential results in Section 5 are also consistent. Hence, our modular addition involving $k$ inputs function class is a good data model to analyze and test the model learning ability, i.e., approximation, optimization, and generalization.
@@
High Order Correlation Attention. [76, 3, 4] state that, when $k=3$, $a_{1}+a_{2}+a_{3}\bmod p$ is hard to be captured by traditional attention. Thus, they introduce high-order attention to capture high-order correlation from the input sequence. However, in Section 5, we show that one-layer transformers have a strong learning ability and can successfully learn modular arithmetic tasks even when $k=5$. This implies that the traditional attention may be more powerful than we expect.
@@ Acknowledgement
In Section A, we provide the potential limitations of this work. In Section B, we discuss the societal impacts of our work. In Section C, we provide more related works. In Section D, we introduce some definitions that will be used in the proof. In Section E, we introduce some auxiliary lemma from previous work that we need. In Section F, Section G, Section H, Section I, we provide the proof of our Lemmas and our main results. In particular, we provide two versions of proof (1) $k=3$ and (2) general $k\geq 3$. We use $k=3$ version to illustrate our proof intuition and then extend our proof to the general $k$ version. Finally, in Section J, we provide more experimental results and implementation details.
Appendix A Limitations
Our work has made progress in exploring how neural networks and Transformers can solve complex mathematical problems such as modular addition operation, but the practical application scope of their conclusions is limited.
On the other hand, we admit that our theorem can provide intuition but cannot fully explain the phenomena shown in Figure 4. Thus, we would like to introduce this more general data setting to the community so that we can study and understand grokking in a more broad way. Studying the relationship between the number of neurons and the grokking strength is interesting and important, and we will leave it as our future work.
Appendix B Societal Impact
@@
Appendix C More Related Work
Theoretical Work About Fourier Transform. To calculate Fourier transform there are two main methodologies: one uses carefully chosen samples through hashing functions (referenced in works like [37, 36, 43, 44]) to achieve sublinear sample complexity and running time, while the other uses random samples (as discussed in [10, 35, 65]) with sublinear sample complexity but nearly linear running time. There are many other works studying Fourier transform [79, 39, 32, 56, 18, 81, 16, 82, 19, 88].
Appendix D More Notations and Definitions
@@
Definition D.1 ($L_{b}$ (vector) norm).
@@
Definition D.2 ($L_{a,b}$ (matrix) norm).
@@
Appendix E Tools from Previous Work
Section E.1 states that we can use the single neuron level optimization to get the maximum-margin network. Section E.2 introduces the maximum-margin for multi-class.
E.1 Tools from Previous Work: Implying Single/Combined Neurons
Lemma E.1 (Lemma 5 in page 8 in [58]).
@@ E.1 Tools from Previous Work: Implying Single/Combined Neurons
E.2 Tools from Previous Work: Maximum Margin for Multi-Class
Lemma E.2 (Lemma 6 in page 8 in [58]).
@@ E.2 Tools from Previous Work: Maximum Margin for Multi-Class
Appendix F Class-weighted Max-margin Solution of Single Neuron
Section F.1 introduces some definitions. Section F.2 shows how we transfer the problem to discrete Fourier space. Section F.3 proposes the weighted margin of the single neuron. Section F.4 shows how we transfer the problem to discrete Fourier space for general $k$ version. Section F.5 provides the solution set for general $k$ version and the maximum weighted margin for a single neuron.
F.1 Definitions
Definition F.1.
@@ F.1 Definitions
Definition F.2.
Let $\eta$ be defined in Definition F.1. When $k=3$, provided the following conditions are met
@@
F.2 Transfer to Discrete Fourier Space
@@
  • We define $\Omega_{q}^{{}^{\prime}*}$ in Definition F.2.
@@ F.3 Get Solution Set
  • We define $\Omega_{q}^{{}^{\prime}*}$ in Definition F.2.
@@
By Lemma F.3, we only need to maximize Equation (F.2).
@@
Consider the odd $p$, Equation (F.2) becomes:
@@
where the first step comes from definition (F.2), the second step follows from Eq. (8), the third step comes from $\widehat{u}_{1}(-j)=\overline{\widehat{u}_{1}(j)}$ and $\widehat{u}_{1}(j)=|\widehat{u}_{1}(j)|\exp(\mathbf{i}\theta_{u_{1}}(j))$, the last step follow from Euler’s formula.
@@
F.4 Transfer to Discrete Fourier Space for General $k$ Version
Definition F.5.
@@ F.4 Transfer to Discrete Fourier Space for General $k$ Version
Definition F.6.
Let $\eta$ be defined in Definition F.5. Provided the following conditions are met
@@
  • We define $\Omega_{q}^{{}^{\prime}*}$ in Definition F.6.
@@
F.5 Get Solution Set for General $k$ Version
Lemma F.8 (Formal version of Lemma 4.2).
@@ F.5 Get Solution Set for General $k$ Version
  • Let $\Omega_{q}^{{}^{\prime}*}$ be defined as Definition F.6.
@@
By Lemma F.7, we only need to maximize Equation (F.4). Thus, the mass of $\widehat{u}_{1},\dots,\widehat{u}_{k}$, and $\widehat{w}$ must be concentrated on the same frequencies. For all $j\in\mathbb{Z}_{p}$, we have
@@
Considering odd $p$, Equation (F.4) becomes:
@@
where the first step follows from definition (F.4), the second step comes from Eq. (17), the last step follows from Eq. (18), i.e., Euler’s formula.
@@
Appendix G Construct Max Margin Solution
Section G.1 proposed the sum-to-product identities for $k$ inputs. Section G.2 shows how we construct $\theta^{*}$ when $k=3$. Section G.3 gives the constructions for $\theta^{*}$ for general $k$ version.
G.1 Sum-to-product Identities
Lemma G.1 (Sum-to-product Identities).
@@ G.1 Sum-to-product Identities
G.2 Constructions for $\theta^{*}$
Lemma G.2.
@@ G.2 Constructions for $\theta^{*}$
  • We define $\Omega_{q}^{{}^{\prime}*}$ in Definition F.2.
@@
Referencing Lemma F.4, we can identify elements within $\Omega_{q}^{{}^{\prime}}$. Our set $\theta^{*}$ will be composed of $16(p-1)$ neurons, including $32$ neurons dedicated to each frequency in the range $1,\ldots,\frac{p-1}{2}$. Focusing on a specific frequency $\zeta$, for the sake of simplicity, let us use $\cos_{\zeta}(x)$ to represent $\cos(2\pi\zeta x/p)$ and $\sin_{\zeta}(x)$ likewise. We note:
@@
Similarly, all other six terms in Eq. (G.2) can be composed by four neurons with different $(\theta_{u_{1}}^{*},\theta_{u_{2}}^{*},\theta_{u_{3}}^{*},\theta_{w}^{*})$.
@@
G.3 Constructions for $\theta^{*}$ for General $k$ Version
Lemma G.3 (Formal version of Lemma 4.3).
@@ G.3 Constructions for $\theta^{*}$ for General $k$ Version
  • We define $\Omega_{q}^{{}^{\prime}*}$ in Definition F.2.
@@
By Lemma F.4, we can get elements of $\Omega_{q}^{{}^{\prime}*}$. Our set $\theta^{*}$ will be composed of $2^{2k-1}\cdot{p-1\over 2}$ neurons, including $2^{2k-1}$ neurons dedicated to each frequency in the range $1,\ldots,\frac{p-1}{2}$. Focusing on a specific frequency $\zeta$, for the sake of simplicity, let us use $\cos_{\zeta}(x)$ to represent $\cos(2\pi\zeta x/p)$ and $\sin_{\zeta}(x)$ likewise.
@@
Appendix H Check Fourier Frequencies
Section H.1 proves all frequencies are used. Section H.2 proves all frequencies are used for general $k$ version.
H.1 All Frequencies are Used
@@
In this proof, let $j_{1},j_{2},j_{3},j_{4}\in\mathbb{Z}$, and $\theta_{u}=\theta_{u}^{*}\cdot\frac{p}{2\pi}$ to simplify the notation. By Lemma F.4,
@@
Setting aside the previously discussed points, it’s established in Lemma E.2 that the function $f$ maintains a consistent margin for various inputs as well as over different classes, i.e., $f$ can be broken down as
@@
H.2 All Frequencies are Used for General $k$ Version

### v4

@@ 5.3 Grokking under Different $k$
To support the importance of our data setting, we study the grokking phenomenon in our data distribution. Following the experiments’ protocol in [76], we show there is the grokking phenomenon under different $k$. We train two-layer transformers with $m=160$ attention heads and hidden dimension as 128 on $k=2,3,4,5$-sum mod-$p=97,31,11,5$ addition dataset with 50% of the data in training. We use different $p$ to guarantee the dataset sizes are roughly equal to each other. Figure 4 shows that the grokking weakens as the number of $k$ increases, which is consistent with our analysis. When $k$ increases, the function class will become more complicated, as we may need more neurons to achieve the max-margin solution. Thus, we use our Theorem 4.1 as a metric to measure the data complexity. It implies that when the ground-truth function class becomes “complicated”, the transformers need to train more steps to fit the training datasets, and the generalization tends to be better. Brilliant recent works by [51, 46] argue that, during learning, the network will be first in the lazy training/NTK regime and then transfer to the rich/feature learning regime sharply, leading to a grokking phenomenon. We use learning steps required for regime switch as a metric of grokking strength.
@@
“Underfitting” in NTK but “overfitting” in Feature Learning. NTK is a notorious overparameterized regime, which probably needs a much larger number of neurons than our max-margin convergence case, i.e., much larger than $\Omega(2^{2k})$ in Theorem 4.1. Thus, under the fixed $m$ and increasing $k$, the model may easily escape the NTK regime, or there is no longer an NTK regime. Thus, we will see a weaker grokking phenomenon as the learning steps needed to transfer from the NTK regime to the feature learning regime become fewer. With increasing $k$, the model will have an “underfitting” issue in the NTK regime, meaning the model must need feature learning to fit the task but cannot only fit the task by NTK. However, the model still has an “overfitting” in the feature learning regime.
@@ 6.1 Grokking in Transformers
The interpretability of grokking in Transformers is explored in [70]. By examining various intermediate states within the residual stream of the Transformer model, it is validated that the model employs Fourier features to tackle the modular addition task. However, fully comprehending how the Transformer model and LLMs perform modular addition remains challenging based on the current work, particularly from a theoretical standpoint. We contend that beginning with a simplistic model setup and achieving a thorough and theoretical understanding of how the network utilizes Fourier features to address the problem serves as a valuable starting point and it provides a theoretical understanding of the grokking phenomenon. We believe that further study on Transformers will be an interesting and important future direction.
6.2 Grokking, Benign Overfitting, and Implicit Bias
@@
Recently, [111] connects the grokking phenomenon to benign overfitting [9, 13, 98, 26, 28]. It shows how the network undergoes a grokking period from catastrophic to benign overfitting. [51, 46] uses implicit bias [83, 30, 42, 91, 69, 12, 58, 39, 109, 110] to explain grokking, where grokking happens if the early phase bias implies an overfitting solution while late phase bias implies a generalizable solution. The intuition from the benign overfitting and the implicit bias well align with our observation in Section 5. It is interesting and valuable to rigorously analyze the grokking or emergent ability under different function class complexities, e.g., Eq (1). We leave this challenge problem as a future work.
6.3 High Order Correlation Attention
@@
[84, 3, 4, 61, 55, 114] state that, when $k=3$, $a_{1}+a_{2}+a_{3}\bmod p$ is hard to be captured by traditional attention. Thus, they introduce high-order attention to capture high-order correlation from the input sequence. However, in Section 5, we show that one-layer transformers have a strong learning ability and can successfully learn modular arithmetic tasks even when $k=5$. This implies that the traditional attention may be more powerful than we expect.
6.4 Connection to Parity and SQ Hardness
If we let $p=2$, then $(a_{1}+\dots+a_{k})\bmod p$ will degenerate to parity function, i.e., $b_{1},\dots,b_{k}\in\{\pm 1\}$ and determining $\prod_{i=1}^{k}b_{i}$. Parity functions serve as a fundamental set of learning challenges in computational learning theory, often used to demonstrate computational obstacles [88]. In particular, $(n,k)$-sparse parity problem is notorious hard to learn, i.e., Statistical Query (SQ) hardness [8]. [23] showed that one-hidden layer networks need an $\Omega(\exp(k))$ number of neurons or an $\Omega(\exp(k))$ number of training steps to successfully learn it by SGD. In our work, we are studying Eq. (2), which is a more general function than parity and indeed is a learning hardness. Our Theorem 4.1 states that we need $\Omega(\exp(k))$ number of neurons to represent the maximum-margin solution, which well aligns with existing works. Our experiential results in Section 5 are also consistent. Hence, our modular addition involving $k$ inputs function class is a good data model to analyze and test the model learning ability, i.e., approximation, optimization, and generalization.
@@ 7 CONCLUSION
8 LIMITATIONS
Our work has made progress in exploring how neural networks and Transformers can solve complex mathematical problems such as modular addition operation, but the practical application scope of their conclusions is limited. On the other hand, we admit that our theorem can provide intuition but cannot fully explain the phenomena shown in Figure 4. Thus, we would like to introduce this more general data setting to the community so that we can study and understand grokking in a more broad way. Studying the relationship between the number of neurons and the grokking strength is interesting and important, and we will leave it as our future work.
@@ Acknowledgement
In Section A, we discuss the societal impacts of our work. In Section B, we introduce some definitions that will be used in the proof. In Section C, we introduce some auxiliary lemma from previous work that we need. In Section D, Section E, Section F, Section G, we provide the proof of our Lemmas and our main results. In particular, we provide two versions of proof (1) $k=3$ and (2) general $k\geq 3$. We use $k=3$ version to illustrate our proof intuition and then extend our proof to the general $k$ version. Finally, in Section H, we provide more experimental results and implementation details.
Appendix A Societal Impact
@@
Appendix B More Notations and Definitions
@@
Definition B.1 ($L_{b}$ (vector) norm).
@@
Definition B.2 ($L_{a,b}$ (matrix) norm).
@@
Appendix C Tools from Previous Work
Section C.1 states that we can use the single neuron level optimization to get the maximum-margin network. Section C.2 introduces the maximum-margin for multi-class.
C.1 Tools from Previous Work: Implying Single/Combined Neurons
Lemma C.1 (Lemma 5 in page 8 in [65]).
@@ C.1 Tools from Previous Work: Implying Single/Combined Neurons
C.2 Tools from Previous Work: Maximum Margin for Multi-Class
Lemma C.2 (Lemma 6 in page 8 in [65]).
@@ C.2 Tools from Previous Work: Maximum Margin for Multi-Class
Condition C.3 (Condition C.1 in page 8 in [65]).
Appendix D Class-weighted Max-margin Solution of Single Neuron
Section D.1 introduces some definitions. Section D.2 shows how we transfer the problem to discrete Fourier space. Section D.3 proposes the weighted margin of the single neuron. Section D.4 shows how we transfer the problem to discrete Fourier space for general $k$ version. Section D.5 provides the solution set for general $k$ version and the maximum weighted margin for a single neuron.
D.1 Definitions
Definition D.1.
@@ D.1 Definitions
Definition D.2.
Let $\eta$ be defined in Definition D.1. When $k=3$, provided the following conditions are met
@@
D.2 Transfer to Discrete Fourier Space
@@
  • We define $\Omega_{q}^{{}^{\prime}*}$ in Definition D.2.
@@ D.3 Get Solution Set
  • We define $\Omega_{q}^{{}^{\prime}*}$ in Definition D.2.
@@
By Lemma D.3, we only need to maximize Equation (D.2).
@@
Consider the odd $p$, Equation (D.2) becomes:
@@
where the first step comes from definition (D.2), the second step follows from Eq. (9), the third step comes from $\widehat{u}_{1}(-j)=\overline{\widehat{u}_{1}(j)}$ and $\widehat{u}_{1}(j)=|\widehat{u}_{1}(j)|\exp(\mathbf{i}\theta_{u_{1}}(j))$, the last step follow from Euler’s formula.
@@
D.4 Transfer to Discrete Fourier Space for General $k$ Version
Definition D.5.
@@ D.4 Transfer to Discrete Fourier Space for General $k$ Version
Definition D.6.
Let $\eta$ be defined in Definition D.5. Provided the following conditions are met
@@
  • We define $\Omega_{q}^{{}^{\prime}*}$ in Definition D.6.
@@
D.5 Get Solution Set for General $k$ Version
Lemma D.8 (Formal version of Lemma 4.2).
@@ D.5 Get Solution Set for General $k$ Version
  • Let $\Omega_{q}^{{}^{\prime}*}$ be defined as Definition D.6.
@@
By Lemma D.7, we only need to maximize Equation (D.4). Thus, the mass of $\widehat{u}_{1},\dots,\widehat{u}_{k}$, and $\widehat{w}$ must be concentrated on the same frequencies. For all $j\in\mathbb{Z}_{p}$, we have
@@
Considering odd $p$, Equation (D.4) becomes:
@@
where the first step follows from definition (D.4), the second step comes from Eq. (18), the last step follows from Eq. (19), i.e., Euler’s formula.
@@
Appendix E Construct Max Margin Solution
Section E.1 proposed the sum-to-product identities for $k$ inputs. Section E.2 shows how we construct $\theta^{*}$ when $k=3$. Section E.3 gives the constructions for $\theta^{*}$ for general $k$ version.
E.1 Sum-to-product Identities
Lemma E.1 (Sum-to-product Identities).
@@ E.1 Sum-to-product Identities
E.2 Constructions for $\theta^{*}$
Lemma E.2.
@@ E.2 Constructions for $\theta^{*}$
  • We define $\Omega_{q}^{{}^{\prime}*}$ in Definition D.2.
@@
Referencing Lemma D.4, we can identify elements within $\Omega_{q}^{{}^{\prime}}$. Our set $\theta^{*}$ will be composed of $16(p-1)$ neurons, including $32$ neurons dedicated to each frequency in the range $1,\ldots,\frac{p-1}{2}$. Focusing on a specific frequency $\zeta$, for the sake of simplicity, let us use $\cos_{\zeta}(x)$ to represent $\cos(2\pi\zeta x/p)$ and $\sin_{\zeta}(x)$ likewise. We note:
@@
Similarly, all other six terms in Eq. (E.2) can be composed by four neurons with different $(\theta_{u_{1}}^{*},\theta_{u_{2}}^{*},\theta_{u_{3}}^{*},\theta_{w}^{*})$.
@@
E.3 Constructions for $\theta^{*}$ for General $k$ Version
Lemma E.3 (Formal version of Lemma 4.3).
@@ E.3 Constructions for $\theta^{*}$ for General $k$ Version
  • We define $\Omega_{q}^{{}^{\prime}*}$ in Definition D.2.
@@
By Lemma D.4, we can get elements of $\Omega_{q}^{{}^{\prime}*}$. Our set $\theta^{*}$ will be composed of $2^{2k-1}\cdot{p-1\over 2}$ neurons, including $2^{2k-1}$ neurons dedicated to each frequency in the range $1,\ldots,\frac{p-1}{2}$. Focusing on a specific frequency $\zeta$, for the sake of simplicity, let us use $\cos_{\zeta}(x)$ to represent $\cos(2\pi\zeta x/p)$ and $\sin_{\zeta}(x)$ likewise.
@@
Appendix F Check Fourier Frequencies
Section F.1 proves all frequencies are used. Section F.2 proves all frequencies are used for general $k$ version.
F.1 All Frequencies are Used
@@
In this proof, let $j_{1},j_{2},j_{3},j_{4}\in\mathbb{Z}$, and $\theta_{u}=\theta_{u}^{*}\cdot\frac{p}{2\pi}$ to simplify the notation. By Lemma D.4,
@@
Setting aside the previously discussed points, it’s established in Lemma C.2 that the function $f$ maintains a consistent margin for various inputs as well as over different classes, i.e., $f$ can be broken down as
@@
F.2 All Frequencies are Used for General $k$ Version